OpenStudy (anonymous):

solve the system x+4y-z=3 (1) x+5y-2z=5 (2) 3x+12y-3z=9 (3)

OpenStudy (lgbasallote):

how did you start it?

OpenStudy (anonymous):

substitution by elimination. x-4y+2z=3 (1), x+5y-2z=5 (2)

OpenStudy (anonymous):

Multiply both the sides by (-1)..

OpenStudy (anonymous):

so x=-1-3y?

OpenStudy (anonymous):

\(x + 3y =-1\)

OpenStudy (anonymous):

cool. thanks

OpenStudy (anonymous):

Welcome dear..

OpenStudy (turingtest):

?

OpenStudy (turingtest):

um... nobody solved anything

OpenStudy (anonymous):

What is the meaning??

OpenStudy (lgbasallote):

lol

OpenStudy (turingtest):

you are only give two equations?

OpenStudy (turingtest):

...and 3 variables?

OpenStudy (anonymous):

Question is not that..

OpenStudy (lgbasallote):

and i dont know where -x - 3y = 1 came either...maybe that's the third

OpenStudy (turingtest):

@jrzyby1 please type the original question in full

OpenStudy (anonymous):

hether or not i need to multil ply by -1 or not

OpenStudy (anonymous):

I think he is just looking for how to write the given equation in standard form.. Coefficient of x is -1 so we can make it one..

OpenStudy (anonymous):

correct

OpenStudy (turingtest):

@jrzyby1 what is this: substitution by elimination. x-4y+2z=3 (1), x+5y-2z=5 (2) you typed this and said that is where you got what you asked us about from

OpenStudy (lgbasallote):

but still where did -x - 3y = 1 came from

OpenStudy (turingtest):

that too

OpenStudy (lgbasallote):

and x + 3y = -1 isnt standard form either...too many questions o.O

OpenStudy (anonymous):

He asked it directly and randomly you can say that..

OpenStudy (anonymous):

yes. that was what I got and didn't know how to finish the equation I wanted to find out if multiplying the last equation by (-10 would give my the variable

OpenStudy (anonymous):

oops (-1)

OpenStudy (lgbasallote):

give you the variable?

OpenStudy (anonymous):

Now I am in doubt too.. Ha ha ha..

OpenStudy (turingtest):

@jrzyby1 are you solving a system of equations? x-4y+2z=3 (1), x+5y-2z=5 (2) if not, why did you type this?

OpenStudy (turingtest):

and I didn't hear anybody ask for standard form either

OpenStudy (anonymous):

I was sustituting-2 into the first equation and added it to the second which gave me the result in question. I didn't know if I needed to multiply it by (-1) to finish it.

OpenStudy (turingtest):

so in the original question, you are given only 2 equations with 3 variables correct?

OpenStudy (turingtest):

x-4y+2z=3 (1) x+5y-2z=5 (2) is this the \(original\) problem?

OpenStudy (lgbasallote):

if you multiplied the first equation by 2..then where's the z variable?

OpenStudy (anonymous):

oh my, the 3 equation ended in zeros. 3x+12y-3z=9

OpenStudy (turingtest):

@jrzyby1 so the original problem is x-4y+2z=3 (1) x+5y-2z=5 (2) 3x+12y-3z=9 (3) correct?

OpenStudy (anonymous):

correct

OpenStudy (turingtest):

now we are getting somewhere...

OpenStudy (turingtest):

please type the problem in full next time, you have not done it correctly so far and have not provided enough info to solve it even if you had

OpenStudy (turingtest):

I am going to modify your post for you so you can get a correct answer the one you have above is wrong, I promise you is that okay with you?

OpenStudy (anonymous):

wait your 4 in the (1) is (+)

OpenStudy (turingtest):

x+4y+2z=3 (1) x+5y-2z=5 (2) 3x+12y-3z=9 (3) correct (please double-check)

OpenStudy (anonymous):

hang on. let me giv eyou it again

OpenStudy (turingtest):

yes, please be 100% sure of the original problem

OpenStudy (anonymous):

x+4y-z=3(1) x+5y-2z=5(2) 3x+12y-3z+9(3)

OpenStudy (turingtest):

okay, I think you should repost this problem in full...

OpenStudy (anonymous):

so am I wrong so far?

OpenStudy (turingtest):

well you still didn't type the problem right... you must be more careful if we are to help you.. is not equation (3) supposed to have an = in it?

OpenStudy (turingtest):

solve the system x+4y-z=3 (1) x+5y-2z=5 (2) 3x+12y-3z=9 (3) is what I am guessing the original problem is

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

I'm going to repost this problem for you, okay? that way you will get the most help I think

OpenStudy (anonymous):

k

OpenStudy (anonymous):

sorry for the hassle

OpenStudy (turingtest):

it's okay check it out I posted your problem for you, it's at the top